# Siyou Li1  Huanan Wu2  Juexi Shao1  Yinghao Ma1  Yujian Gan1 Yihao Luo3  Yuwei Wang4  Dong Nie5  Lu Wang6  Wengqing Wu1,7 Le Zhang1,8  Massimo Poesio1,9  Juntao Yu1

### v2

@@ 1 Introduction
Recent multimodal large language models (MLLMs) build on strong vision–language pretraining [radford2021learningtransferablevisualmodels] that extend their ability to solve image and video understanding tasks. As the ability of MLLMs grows, long video understanding becomes one of the new frontiers of multimodal AI with benchmarks like Video-MME [fu2024video], Video-MMMU [hu2025videommmu], LVBench [wang2024lvbench], and MLVU [zhou2024mlvu] that emphasize tasks that require reasoning over extended temporal horizons. Understanding such long videos is crucial for applications ranging from assistive robotics and surgical coaching to copyright compliance and safety monitoring. However, most existing video–language models inherit design assumptions from short-clip recognition: they pre-truncate the input or aggressively down-sample frames, discarding fine-grained temporal cues that matter when a question refers to “the third screw tightened” or “the moment the cake first rises”.
@@ 2 Related Works
Recent MLLMs leverage strong vision–language pre‑training and have driven progress on long‑video benchmarks. Orthogonally, efficiency research explores token pruning and routing, visual token grouping/merging, and dynamic inference. The vision encoder, often a Vision Transformer (ViT[dosovitskiy2021imageworth16x16words]), typically produces a large set of patch tokens, which creates a prohibitive computational bottleneck for the LLM’s self-attention mechanisms. To address this, several strategies for visual feature selection and compression have been proposed.
A dominant paradigm is Query-Based Resampling, which acts as an information bottleneck. This method employs a small, fixed-size set of learnable queries that distil information from the dense visual token set via cross-attention. This produces a compact, fixed-length sequence of visual embeddings that summarize the most salient features. Prominent examples of this approach include the Q-Former architecture utilized in BLIP-2[li2023blip2bootstrappinglanguageimagepretraining] and the Perceiver Resampler found in Flamingo[alayrac2022flamingovisuallanguagemodel].
Alternative strategies operate directly on the visual token sequence. Token Pruning methods aim to reduce sequence length by identifying and discarding tokens based on saliency, redundancy, or task-relevance metrics (such as SAINT[jeddi2025similarityawaretokenpruningvlm]). In contrast, Token Merging techniques dynamically fuse similar or spatially adjacent tokens into unified representations, progressively shortening the sequence while attempting to preserve information content. These merging mechanisms are all fundamentally designed to reduce the input sequence length for the LLM, thereby mitigating the quadratic computational cost while retaining the most critical visual information necessary for multimodal reasoning.
Our approach is motivated by the same aim to reduce the number of visual tokens, but additionally takes into account the input query and only retains visual tokens that are highly relevant to the query in question. As a result, our method can reduce the number of visual tokens more significantly than canonical token-merging methods like the ones used in ToMe[bolya2023tokenmergingvitfaster] and Qwen2.5-VL[bai2025qwen25vltechnicalreport]. Take the latest Qwen2.5-VL model as an example. It uses a simple MLP-based merging strategy that merges spatially adjacent tokens (e.g., n=4) into a single representation. As a result, the actual number of tokens fed into the LLM is reduced by a factor of n. Such an approach extends the model’s ability from short videos to medium-length videos. However, when the input video frame rate is excessively high, the vision embedding still becomes extremely large. For example, when a 480p video input reaches 450 frames (using Qwen2.5-VL), the embedding sequence length fed to the language model reaches 136,035. This exceeds the maximum processing length of the model (131,072). In contrast, our approach can reduce the number of visual tokens by up to 89% when compared to the vanilla Qwen2.5-VL model.
@@ 3.2.1 Cross‑Attention Scoring
After the vision head processed the video features, we have a very long sequence of visual features. Meanwhile, the natural‑language question is embedded into a much shorter sequence of text tokens. To decide which visual tokens matter for this particular question, we place a multi‑head cross‑attention layer [vaswani2017attention] between the two streams. Let $\mathcal{Q}{=}[q_{1},\dots,q_{L}]^{\top}\in\mathbb{R}^{L\times d}$ be the text tokens. We compute a cross‑attention map $\alpha\in\mathbb{R}^{h\times L\times M}$ using a Qwen2.5‑compatible attention since we use Qwen2.5-VL as our base model. The per‑token relevance is defined as $r_{i}\;=\;\max_{h,\;\ell}\;\alpha_{h,\ell,i}\;\in[0,1]$ with $r\in[0,1]^{M}$. Intuitively, a visual token that is heavily attended by any word in the question will get a high $r$; tokens the model never looks at stay close to zero.
@@ 3.2.3 Top-n Gate
During training time, we make the gate differentiable to allow gradient flow freely. We choose the threshold $t$ so that the expected number of kept tokens matches the target budget by solving $\sum_{i}\sigma\!\big((r_{i}-t)/\tau_{s}\big)=\rho M$ using a few Newton iterations initialised at the median (Alg. 2). Given $t$, we form per‑token binary keep/drop decisions via a Gumbel‑Softmax [jang2017categorical] straight‑through one‑hot sample $\text{logits}_{i}=\big[r_{i}{-}t,\;0\big],y_{i}=\mathrm{GumbelSoftmax}(\text{logits}_{i};\tau{=}\tau_{s},\;\text{hard}{=}\texttt{true}),s_{\text{keep},i}=y_{i,0}\in\{0,1\}$ with a safety fallback that forces $\max_{i}s_{\text{keep},i}=1$ if all zeros occur. The kept sequence $Z$ preserves the original temporal order.
@@ 4.1 Training Datasets
We train on the ShareGPTVideo [zhang2024direct] corpus containing approximately 300k video–caption pairs. Each video is pre-segmented into frames and stored in a per-video folder to support streaming and batched loading.
@@
Fig. 3 summarizes the pipeline used to transform caption supervision into training signals that are easier for a compact student to learn from. First, for every caption we prompt Qwen3‑235B [qwen3technicalreport] to write a visual single‑choice question with options and a unique correct answer, yielding QTS‑VSCQ1, a VQA dataset with over 855k multiple choices questions. Next, Qwen2.5‑VL acts as an automatic judge: it answers each VSCQ item from the paired video and we keep only items where the model’s prediction matches the provided answer key, forming QTS‑VSCQ2. Finally, we ask Qwen2.5‑VL to produce short free‑form answers (given the video and question), which we pair with the inputs to obtain the QTS‑VQA set for generative supervision.
@@ 4.3 Evaluations on Overheads
Figure 4: Scaling characteristics on a single A100 GPU (FlashAttention[dao2022flashattentionfastmemoryefficientexact] enabled). Left: number of vision embeddings forwarded to the LM versus frame count. Right: end‑to‑end latency to first complete output versus frame count. QTSplus‑3B‑FT consistently reduces the embedding budget and latency relative to Qwen2.5‑VL‑3B‑Instruct.
@@ 4.4.1 Framework
We use the popular open-source multimodal evaluation framework lmms-eval[zhang2024lmmsevalrealitycheckevaluation]. This framework aggregates many popular multimodal benchmarks and provides most datasets, enabling efficient multi-dataset, multi-model testing. The following results are based on datasets and subsets integrated by the framework. To support inference for multimodal LMs using QTSplus, we made minor modifications to the framework, including some bug fixes and adaptation for QTSplus models.
Some datasets stored in the lmms-eval repository may differ from their official versions; for compatibility, certain implementation details may also vary. We observed feedback that results obtained with this tool can diverge from leaderboard numbers. We believe that the underlying LLM inference environment evolves over time, making it hard to exactly match results from a year ago. For example, some test suites (e.g., TempoCompass[chen2025mega-bench]) use GPT-3.5 Turbo for assisted verification. Since this service is inaccessible in our region, we use the latest QWen3-235B-A22B-Instruct as a substitute. As these LLMs are only used to assess answer alignment, the impact on results should be limited.
@@ 4.5 Evaluations on Open Datasets
TempoCompass [chen2025mega-bench] checks the fine grained understanding of videos. Video‑MME [fu2024video] measures broad video understanding across scenarios. Video‑MMMU [hu2025videommmu] targets diverse, expert‑level tasks and reports adaptation/perception/comprehension splits (we follow the adaptation metric). LVBench [wang2024lvbench] emphasizes long videos up to two hours. MLVU [zhou2024mlvu] focuses on multimodal long‑video understanding with global and local reasoning. MVBench [li2024mvbenchcomprehensivemultimodalvideo] contains 20 temporal sub‑tasks (e.g., action count/sequence, object interaction, scene transition).
@@ Appendix C Correlation Analysis between predicted budget $\rho$ and gating threshold $t$
Table 3: Performance of Different Components
Datasets UNIF nREENC QTS3B
General Multimodal Tasks
TCaction_accuracy 94.67 95.56 96.75
TCattribute_change_acc 57.99 60.76 67.01
TCavg_accuracy 59.94 61.08 64.18
TCdirection_accuracy 38.21 39.10 39.10
TCorder_accuracy 58.61 61.26 67.88
TCspeed_accuracy 48.90 47.63 49.84
Video Understanding Tasks
Video-MME 47.63 51.48 53.56
LVBench 32.92 33.38 36.86
MLVUtest 13.45 9.82 22.51
VMadaptation 25.33 25.33 25.67
VMcomprehension 34.67 33.67 35.67
VMperception 39.33 40.67 47.00
MBaction_antonym 79.00 79.50 80.00
MBaction_count 53.50 53.50 62.00
MBaction_localization 41.50 44.00 42.00
MBaction_prediction 57.50 59.50 63.50
MBaction_sequence 66.50 65.00 67.50
MBcharacter_order 56.00 68.50 71.50
MBcounterfactual_inference 56.50 67.00 74.00
MBegocentric_navigation 39.00 39.50 41.00
MBepisodic_reasoning 49.50 51.50 56.50
MBfine_grained_action 47.00 47.00 46.50
MBfine_grained_pose 50.50 54.50 60.00
MBmoving_attribute 75.00 85.50 89.50
MBmoving_count 56.00 60.00 65.00
MBmoving_direction 37.00 40.50 41.50
MBobject_existence 74.50 85.00 86.00
MBobject_interaction 66.00 68.00 67.50
MBobject_shuffle 32.50 33.50 34.00
MBscene_transition 89.00 88.00 89.50
MBstate_change 53.00 53.50 55.00
MBunexpected_action 76.00 81.00 82.00
Appendix D Ablation Study
Protocol and variants
  1. 1.
    UNIF: replace the whole QTSplus layer with uniform sampling.
  2. 2.
  3. 3.
    QTS3B: full configuration (query-aware selector ++ re-encoder).
Aggregate effects
  • Video-MME: +5.93; +2.08.
  • LVBench: +3.94; +3.48.
  • MLVUtest: +9.06; +12.69.
Fine-grained MVBench analysis.
Summary
@@ Appendix D Ablation Study
Appendix E Training hyperparameters

### v3

# Siyou Li1  Huanan Wu2  Juexi Shao1  Yinghao Ma1  Yujian Gan1 Yihao Luo3  Yuwei Wang4  Dong Nie5  Lu Wang6  Wenqing Wu1,7 Le Zhang1,8  Massimo Poesio1,9  Juntao Yu1

@@ 1 Introduction
Recent multimodal large language models (MLLMs) build on strong vision–language pretraining [13] that extend their ability to solve image and video understanding tasks. As the ability of MLLMs grows, long video understanding becomes one of the new frontiers of multimodal AI with benchmarks like Video-MME [7], Video-MMMU [8], LVBench [16], and MLVU [19] that emphasize tasks that require reasoning over extended temporal horizons. Understanding such long videos is crucial for applications ranging from assistive robotics and surgical coaching to copyright compliance and safety monitoring. However, most existing video–language models inherit design assumptions from short-clip recognition: they pre-truncate the input or aggressively down-sample frames, discarding fine-grained temporal cues that matter when a question refers to “the third screw tightened” or “the moment the cake first rises”.
@@ 2 Related Works
Recent MLLMs leverage strong vision–language pre‑training and have driven progress on long‑video benchmarks. Orthogonally, efficiency research explores token pruning and routing, visual token grouping/merging, and dynamic inference. The vision encoder, often a Vision Transformer (ViT[6]), typically produces a large set of patch tokens, which creates a prohibitive computational bottleneck for the LLM’s self-attention mechanisms. To address this, several strategies for visual feature selection and compression have been proposed.
A dominant paradigm is Query-Based Resampling, which acts as an information bottleneck. This method employs a small, fixed-size set of learnable queries that distil information from the dense visual token set via cross-attention. This produces a compact, fixed-length sequence of visual embeddings that summarize the most salient features. Prominent examples of this approach include the Q-Former architecture utilized in BLIP-2[11] and the Perceiver Resampler found in Flamingo[1].
Alternative strategies operate directly on the visual token sequence. Token Pruning methods aim to reduce sequence length by identifying and discarding tokens based on saliency, redundancy, or task-relevance metrics (such as SAINT[10]). In contrast, Token Merging techniques dynamically fuse similar or spatially adjacent tokens into unified representations, progressively shortening the sequence while attempting to preserve information content. These merging mechanisms are all fundamentally designed to reduce the input sequence length for the LLM, thereby mitigating the quadratic computational cost while retaining the most critical visual information necessary for multimodal reasoning.
Our approach is motivated by the same aim to reduce the number of visual tokens, but additionally takes into account the input query and only retains visual tokens that are highly relevant to the query in question. As a result, our method can reduce the number of visual tokens more significantly than canonical token-merging methods like the ones used in ToMe[3] and Qwen2.5-VL[2]. Take the latest Qwen2.5-VL model as an example. It uses a simple MLP-based merging strategy that merges spatially adjacent tokens (e.g., n=4) into a single representation. As a result, the actual number of tokens fed into the LLM is reduced by a factor of n. Such an approach extends the model’s ability from short videos to medium-length videos. However, when the input video frame rate is excessively high, the vision embedding still becomes extremely large. For example, when a 480p video input reaches 450 frames (using Qwen2.5-VL), the embedding sequence length fed to the language model reaches 136,035. This exceeds the maximum processing length of the model (131,072). In contrast, our approach can reduce the number of visual tokens by up to 89% when compared to the vanilla Qwen2.5-VL model.
@@ 3.2.1 Cross‑Attention Scoring
After the vision head processed the video features, we have a very long sequence of visual features. Meanwhile, the natural‑language question is embedded into a much shorter sequence of text tokens. To decide which visual tokens matter for this particular question, we place a multi‑head cross‑attention layer [15] between the two streams. Let $\mathcal{Q}{=}[q_{1},\dots,q_{L}]^{\top}\in\mathbb{R}^{L\times d}$ be the text tokens. We compute a cross‑attention map $\alpha\in\mathbb{R}^{h\times L\times M}$ using a Qwen2.5‑compatible attention since we use Qwen2.5-VL as our base model. The per‑token relevance is defined as $r_{i}\;=\;\max_{h,\;\ell}\;\alpha_{h,\ell,i}\;\in[0,1]$ with $r\in[0,1]^{M}$. Intuitively, a visual token that is heavily attended by any word in the question will get a high $r$; tokens the model never looks at stay close to zero.
@@ 3.2.3 Top-n Gate
During training time, we make the gate differentiable to allow gradient flow freely. We choose the threshold $t$ so that the expected number of kept tokens matches the target budget by solving $\sum_{i}\sigma\!\big((r_{i}-t)/\tau_{s}\big)=\rho M$ using a few Newton iterations initialised at the median (Alg. 2). Given $t$, we form per‑token binary keep/drop decisions via a Gumbel‑Softmax [9] straight‑through one‑hot sample $\text{logits}_{i}=\big[r_{i}{-}t,\;0\big],y_{i}=\mathrm{GumbelSoftmax}(\text{logits}_{i};\tau{=}\tau_{s},\;\text{hard}{=}\texttt{true}),s_{\text{keep},i}=y_{i,0}\in\{0,1\}$ with a safety fallback that forces $\max_{i}s_{\text{keep},i}=1$ if all zeros occur. The kept sequence $Z$ preserves the original temporal order.
@@ 4.1 Training Datasets
We train on the ShareGPTVideo [18] corpus containing approximately 300k video–caption pairs. Each video is pre-segmented into frames and stored in a per-video folder to support streaming and batched loading.
@@
Fig. 3 summarizes the pipeline used to transform caption supervision into training signals that are easier for a compact student to learn from. First, for every caption we prompt Qwen3‑235B [14] to write a visual single‑choice question with options and a unique correct answer, yielding QTS‑VSCQ1, a VQA dataset with over 855k multiple choices questions. Next, Qwen2.5‑VL acts as an automatic judge: it answers each VSCQ item from the paired video and we keep only items where the model’s prediction matches the provided answer key, forming QTS‑VSCQ2. Finally, we ask Qwen2.5‑VL to produce short free‑form answers (given the video and question), which we pair with the inputs to obtain the QTS‑VQA set for generative supervision.
@@ 4.3 Evaluations on Overheads
Figure 4: Scaling characteristics on a single A100 GPU (FlashAttention[5] enabled). Left: number of vision embeddings forwarded to the LM versus frame count. Right: end‑to‑end latency to first complete output versus frame count. QTSplus‑3B‑FT consistently reduces the embedding budget and latency relative to Qwen2.5‑VL‑3B‑Instruct.
@@ 4.4.1 Framework
We use the popular open-source multimodal evaluation framework lmms-eval[17]. This framework aggregates many popular multimodal benchmarks and provides most datasets, enabling efficient multi-dataset, multi-model testing. The following results are based on datasets and subsets integrated by the framework. To support inference for multimodal LMs using QTSplus, we made minor modifications to the framework, including some bug fixes and adaptation for QTSplus models.
Some datasets stored in the lmms-eval repository may differ from their official versions; for compatibility, certain implementation details may also vary. We observed feedback that results obtained with this tool can diverge from leaderboard numbers. We believe that the underlying LLM inference environment evolves over time, making it hard to exactly match results from a year ago. For example, some test suites (e.g., TempoCompass[4]) use GPT-3.5 Turbo for assisted verification. Since this service is inaccessible in our region, we use the latest QWen3-235B-A22B-Instruct as a substitute. As these LLMs are only used to assess answer alignment, the impact on results should be limited.
@@ 4.5 Evaluations on Open Datasets
TempoCompass [4] checks the fine grained understanding of videos. Video‑MME [7] measures broad video understanding across scenarios. Video‑MMMU [8] targets diverse, expert‑level tasks and reports adaptation/perception/comprehension splits (we follow the adaptation metric). LVBench [16] emphasizes long videos up to two hours. MLVU [19] focuses on multimodal long‑video understanding with global and local reasoning. MVBench [12] contains 20 temporal sub‑tasks (e.g., action count/sequence, object interaction, scene transition).
4.6 Generalization
Model VideoMME MLVU LongVB LVBench
LLaVA-Video-7B 54.11 63.78 49.47 40.03
+ QTSplus 52.89 60.93 13.51* 36.67
InternVL2.5-8B 54.78 64.93 48.53 40.87
+ QTSplus 54.22 60.70 39.75 36.99
Table 2: Generalization of QTSplus to LLaVA and InternVL.
(*) indicates lack of specific format tuning during rebuttal.
To demonstrate that QTSplus is model-agnostic, we extended it to two additional state-of-the-art video-LLMs: LLaVA-Video-7B and InternVL2.5-8B. We applied QTSplus (frozen vision encoder and language model, trainable adapter) without extensive tuning.
As shown in Table 2, QTSplus successfully generalizes. On InternVL2.5, QTSplus retains $\approx$99% of performance on VideoMME and LVBench while reducing the visual token count by $\approx$89%. The drop on LongVB for LLaVA is due to the lack of instruction tuning on that specific format in the limited rebuttal time, yet the high retention on VideoMME proves the module effectively selects informative tokens across different backbones.
Table 3: Performance of Different Components
Datasets UNIF nREENC QTS3B
General Multimodal Tasks
TCaction_accuracy 94.67 95.56 96.75
TCattribute_change_acc 57.99 60.76 67.01
TCavg_accuracy 59.94 61.08 64.18
TCdirection_accuracy 38.21 39.10 39.10
TCorder_accuracy 58.61 61.26 67.88
TCspeed_accuracy 48.90 47.63 49.84
Video Understanding Tasks
Video-MME 47.63 51.48 53.56
LVBench 32.92 33.38 36.86
MLVUtest 13.45 9.82 22.51
VMadaptation 25.33 25.33 25.67
VMcomprehension 34.67 33.67 35.67
VMperception 39.33 40.67 47.00
MBaction_antonym 79.00 79.50 80.00
MBaction_count 53.50 53.50 62.00
MBaction_localization 41.50 44.00 42.00
MBaction_prediction 57.50 59.50 63.50
MBaction_sequence 66.50 65.00 67.50
MBcharacter_order 56.00 68.50 71.50
MBcounterfactual_inference 56.50 67.00 74.00
MBegocentric_navigation 39.00 39.50 41.00
MBepisodic_reasoning 49.50 51.50 56.50
MBfine_grained_action 47.00 47.00 46.50
MBfine_grained_pose 50.50 54.50 60.00
MBmoving_attribute 75.00 85.50 89.50
MBmoving_count 56.00 60.00 65.00
MBmoving_direction 37.00 40.50 41.50
MBobject_existence 74.50 85.00 86.00
MBobject_interaction 66.00 68.00 67.50
MBobject_shuffle 32.50 33.50 34.00
MBscene_transition 89.00 88.00 89.50
MBstate_change 53.00 53.50 55.00
MBunexpected_action 76.00 81.00 82.00
5 Ablation Study
Protocol and variants
  1. 1.
    UNIF: replace the whole QTSplus layer with uniform sampling.
  2. 2.
  3. 3.
    QTS3B: full configuration (query-aware selector ++ re-encoder).
Aggregate effects
  • Video-MME: +5.93; +2.08.
  • LVBench: +3.94; +3.48.
  • MLVUtest: +9.06; +12.69.
Fine-grained MVBench analysis.
Summary
@@ Appendix C Correlation Analysis between predicted budget $\rho$ and gating threshold $t$
Appendix D Training hyperparameters